# StoX-Net: Stochastic Processing of Partial Sums for Efficient In-Memory Computing DNN Accelerators

### v1

@@ Abstract.
Crossbar-based in-memory computing (IMC) has emerged as a promising platform for hardware acceleration of deep neural networks (DNNs). However, the energy and latency of IMC systems are dominated by the large overhead of the peripheral analog-to-digital converters (ADCs). To address such ADC bottleneck, here we propose to implement stochastic processing of array-level partial sums (PS) for efficient IMC. Leveraging the probabilistic switching of spin-orbit torque magnetic tunnel junctions, the proposed PS processing eliminates the costly ADC, achieving significant improvement in energy and area efficiency. To mitigate accuracy loss, we develop PS-quantization-aware training that enables backward propagation across stochastic PS. Furthermore, a novel scheme with an inhomogeneous sampling length of the stochastic conversion is proposed. When running ResNet20 on the CIFAR-10 dataset, our architecture-to-algorithm co-design demonstrates up to 22x, 30x, and 142x improvement in energy, latency, and area, respectively, compared to IMC with standard ADC. Our optimized design configuration using stochastic PS achieved 666x (111x) improvement in Energy-Delay-Product compared to IMC with full precision ADC (sparse low-bit ADC), while maintaining near-software accuracy at various benchmark classification tasks.
1. Introduction
@@
Crossbar-based analog in-memory computing (IMC) has demonstrated great potential for Deep Neural Network (DNN) acceleration by achieving high efficiency and massive parallelism at processing matrix-vector-multiplications (MVMs)(Shafiee et al., 2016), which dominates most state-of-the-art DNN workloads(Fleischer et al., 2020). However, the hardware cost of large-scale IMC systems is dominated by the significant overhead of peripheral analog-to-digital converters (ADCs), which is required to ensure robust data communication in a large-scale system. ADC consumes over 60-80 of the energy and chip area of IMC hardware (Chakraborty et al., 2020; Shafiee et al., 2016). Due to the large area overhead of high-precision A-D conversion, an ADC must be shared by multiple columns in a crossbar. Such design leads to a severe throughput bottleneck since the processing of the array-level outputs across multiple columns has to be done sequentially.
@@
While various recent works have aimed to mitigate the ADC bottleneck (Yang et al., 2019; Huang et al., 2021; Saxena et al., 2022), minimizing ADC overhead in IMC while still maintaining satisfactory inference accuracy remains a significant challenge. Various factors contribute to the difficulty in addressing the ADC bottleneck. (1) First, since the array-level partial sums (PS) are not represented at the application level, the standard quantization approach focusing on activation and weights can not directly address the ADC precision for PS. (2) Second, while quantization-aware training is effective at enabling hardware-friendly models, training with PS quantization desires backward propagation with careful incorporation of array-level variables, which requires re-designing the computational graph of the DNN model. (3) Third, the required ADC bit precision for MVM may vary significantly depending on both the algorithmic attributes (such as sparsity and DNN layer dimension) and hardware attributes (such as array size and bits per memory cell). Accommodating these varying scenarios requires reconfigurability and flexibility in ADC, leading to increased overhead and design complexity.
@@
In this work, we propose stochastic processing of the array-level partial sum for efficient IMC leveraging a spin-orbit-torque magnetic tunnel junction (SOT-MTJ) with simple circuitry such as an inverter. To address the accuracy drop due to loss of information at the stochastic PS, a PS-aware training methodology with the stochastic switching behavior of SOT-MTJs is developed. The low-overhead crossbar peripheral based on the proposed spintronic devices/circuits demonstrates significant improvement in energy efficiency and considerable area savings. Our proposed IMC design eliminates the ADC bottleneck and opens up an exciting direction where stochastic computation could play a vital role in designing ultra-efficient non-von Neumann ML accelerators. The major contributions of our work are the following:
@@
We develop a comprehensive framework of PS quantization-aware training that addresses the accuracy degradation due to the stochastic conversion. In particular, both the device-level stochastic MTJ switching behavior, and key architectural attributes including bit slicing and array size, are incorporated into backward propagation. Our model with 1-bit stochastic PS achieves (or slightly exceeds) state-of-the-art accuracy at benchmarking image classification tasks.
@@
We identify that the state-of-the-art quantization of convolutional DNN has severe limitations due to the first convolution layer remaining at high precision. The first convolution layer is typically compute-intensive and dominates the total computation of DNN. To address this challenge, we explore layer-wise inhomogeneous sampling numbers to enable aggressive PS quantization on the first convolution layer. Our inhomogeneous stochastic sampling based on a Monte-Carlo-based sensitivity analysis achieves improvement in accuracy with minimal additional overhead, demonstrating an advantageous trade-off.
2. Background
2.1. DNN acceleration with in-memory computing
@@
Analog IMC is being extensively explored for accelerating DNN inference to meet the growing computational demand. With the weight matrices stored in crossbar arrays, the data movement of MVM processing is drastically reduced, alleviating the von Neumann memory bottleneck. To process a large-scale DNN workload, large matrices and high-precision MVMs need to be partitioned into multiple crossbar arrays. To ensure robust data communication, an analog crossbar IMC macro is connected with other components through digital-to-analog-converters (DAC) at the input and ADC at the crossbar output. The resolution of DAC at each row is typically designed at 1 bit for area saving, while high-precision input activation will be converted to bit streams over multiple time steps (bit streaming). As for high-precision weights, due to the technological limitation in the number of bits per cell, a full-precision weight matrix will be partitioned into several slices of sub-arrays (bit slicing).
@@
The required ADC solution N for array-level PS processing is , where Nrow is the number of activated rows; I is input bits per stream; and W is bits per slice. Since the energy, area, and latency overhead of ADC increase significantly with bit precision (Murmann, [n. d.]), ADC becomes the bottleneck of efficiency and performance in IMC design.
2.2. Stochastic spintronics for machine learning
@@
Several emerging NVM technologies, including resistive memory (ReRAM) and magnetic random access memory (MRAM), have been explored for machine learning acceleration. Recent explorations have demonstrated that spintronic devices exhibit sufficient endurance to realize both synaptic weight storage and neural activation. (Kan et al., 2017) Particularly, an SOT-MTJ under an excitation current with varying magnitude will have the magnetization switching probability as a sigmoidal function of the current, which provides a direct emulation of a stochastic neuron (Sengupta et al., 2016) with significantly higher area/energy efficiency compared to CMOS implementations. Moreover, the separate read/write paths of SOT-MTJs offer immense design flexibility by eliminating the constraint of write current density(Sharma et al., 2021). All-spin neuro-synaptic processors based on stochastic switching of SOT-MTJ have been proposed (Srinivasan et al., 2017). We will explore SOT-MTJs for efficient PS processing of the crossbar output.
2.3. Related work on addressing the ADC bottleneck in IMC
@@
Various works have investigated co-designing the IMC hardware architecture and DNN algorithms to reduce the ADC precision. One major thrust focuses on exploiting quantization with hardware-aware training. In (Kim et al., 2018, 2021) binary neural networks exhibit improved efficiency for IMC, but the models are limited to 1-bit representation. Recent works (Saxena et al., 2022; Kim et al., 2020, 2022) implemented bit slicing/streaming to map workloads with multi-bit weights/activations. Aggressively reducing the ADC precision to 1-bit will essentially enable using sense amplifiers. However, 1-bit PS showed sizeable accuracy degradation even after hardware-aware re-training (Kim et al., 2020, 2022). Moreover, it is important to note that the state-of-the-art quantization-aware training models keep the first convolution layer at full precision (Qin et al., 2020; Saxena et al., 2022; Liu et al., 2018). However, the compute-intensive first convolution layer in image processing can dominate the overall computation workloads. Keeping a full-precision first layer severely limits the overall improvement in energy and throughput.
@@
Another major thrust is to exploit sparsity through pruning and re-training to enable lower ADC precision (Kim et al., 2023; Yang et al., 2019). Higher sparsity reduces the range of possible values of MVM output and thus reduces the ADC resolution requirement. As a result, sparsity-aware IMC with reconfigurable ADC resolution can improve energy and latency. However, it is important to note that the area of the peripheral ADC circuitry remains large in order to handle the highest bit precision of the reconfigurable design. Such large area overhead still hinders the hardware parallelism in IMC architecture.
@@
3. IMC with Stochastic Partial Sums
The proposed Stox-Net is built on IMC crossbar arrays where current-voltage converters, column-shared MUXs, and peripheral ADCs are replaced by a row of stochastic SOT-MTJs (Fig. 2). We first present the device/circuit design of our stochastic PS processing components and then implement stochasticity-aware training of quantized DNNs with bit slicing considered (Liu et al., 2018; Qin et al., 2020).
3.1. Stochastic PS processing based on MTJ
The stochastic processing based on the current-driven probabilistic switching of SOT-MTJs is shown in Fig. 2. The device behavior is simulated using a MATLAB-based macro-spin Landau Lifshitz Gilbert (LLG) equation simulator with the SOT current included. Using an MTJ with tunnel magnetoresistance ratio (TMR) of 350-450% (Sharma et al., 2021), a simple voltage divider circuit combined with a CMOS inverter can behave as a stochastic binary analog to digital converter as shown in Fig. 3(a).
The area of the stochastic MTJ converter is obtained based on the layout in Fig. 3(b). The layout is drawn following the -based design rules (mos, 2004), where is half the feature size. Considering a 28 nm feature size, the area can be estimated as A = 32 * 26 = 0.0163 .
The capability of the stochastic MTJ-converter is illustrated in Fig. 4. For simulating the SOT-MTJ, we use a modified version of the STT-SOT-MTJ model developed by the Spintronics Interdisciplinary Center (Spinlib) (Wang et al., 2015). The MTJ-converter features a reference MTJ and a free SOT-MTJ device arranged as a voltage divider. The SOT-MTJ receives pulses of and of each to imitate the crossbar current. The SOT-MTJ switches its state when is supplied and resets with an equivalent negative current. This current modulation alters the SOT-MTJ’s resistance state, affecting the voltage division ratio with the reference resistor. The midpoint voltage of this divider varies due to the SOT-MTJ’s state and is used as the input to an inverter, and the corresponding output has distinct ‘0’ and ‘1’ voltage levels. The sizing of the transistors in the inverter is optimized by minimizing the width to ensure distinct and clear output levels corresponding to the voltage variations from the divider. The MTJ’s resistance is adjusted by varying its dimensions and the thickness of the oxide layer to achieve a parallel low state resistance of and an anti-parallel high state resistance of , with a TMR of 4.4. The SOT-MTJ’s elliptical shape enhances the efficiency and reliability of the switching process by optimizing the magnetic field distribution within the junction. Our simulations reported a write latency of and read latency of , indicating rapid response times suitable for high-frequency applications. Energy consumption was measured at per conversion cycle by calculating the total power consumption for one period. The device operates over a temperature range of -30°C to 300°C, emphasizing the robustness and potential versatility of the MTJ device in various environmental conditions. As Demonstrated in the following sections, the proposed SOT-MTJ-based converter is orders of magnitude more energy and area-efficient than ADC, leading to significant improvement of system-level hardware efficiency compared to standard IMC.
@@ 3.1. Stochastic PS processing based on MTJ
| VDD | 1 V |
| Ref. MTJ resistance | 435 k |
3.2. Hardware-Aware Training of Quantized DNN with Stochastic Partial Sum
We map large DNN workloads into crossbars with finite array size and limited bits per cell with bit slicing from (Shafiee et al., 2016) and input/weight representation from (Saxena et al., 2022). Mapping a convolution layer of kernel size and input channels using a crossbar array of rows results in a number of subarrays defined by . For bit slicing, we consider 1-, 2-, and 4-bit per memory cell in the StoX-Net design. Based on Algorithm 1, a software-level MVM operation with the number of weight and input slices being and , will have array-level PSs to be shift-and-added (). In addition to workload partitioning, StoX-Net also considers SOT-MTJ samplings in its optimization.
Input: max subarray column length
Output: hardware-aware
To process the crossbar output using the SOT-MTJ, the curve of switching probability versus current is emulated by the tanh function:
where is the sensitivity parameter. Increasing will make the tanh curve more step-like, approaching a deterministic sense amplifier. The effective sensitivity can be altered by tuning the range of crossbar current when mapping MVM operations to hardware.
As shown in Fig. 5, the stochastic network is trained to generate a broader distribution, covering more non-ternary values, which makes leveraging the stochastic switching possible. The deterministic 1-bit sense amplifier (SA) model trains to generate a narrower Gaussian distribution around 0 and a concentration at -1 and 1. During the MTJ conversion, the accumulated analog current supplied to the heavy metal layer of SOT-MTJ will be converted to bi-stable states, which are interpreted as (-1, 1) for better representation in training (Zhu et al., 2020).
To mitigate the errors due to stochastic bit generation, we leverage multisampling to recover accuracy with some trade-off of hardware efficiency. More SOT-MTJ samples will lead to a better representation of the analog output. We explored no more than 8 samples per conversion since the total conversion energy and latency increase linearly with the number of samples.
Additionally, we perform a Monte Carlo simulation on our network’s trainable StoX layers to guide an inhomogeneous multisampling scheme. For each layer, we apply a uniform random perturbation to its weights at inference and then measure the significance of the weights of that layer to the model performance, which is represented by the accuracy loss. Fig. 6 shows the first layer being most susceptible to perturbation, indicating the most significance, while layers close to the output classifications are moderately susceptible to error as well. Based on this analysis, we implement a model with mixed layer-wise sampling numbers (”Mix-QF”), where layers with higher sensitivity utilize more samples and layers with lower sensitivity fewer samples. We achieve an accuracy similar to a homogeneous 4-sample network with only a small increase in operations compared to the 1-sample network. For maintaining higher accuracy and decreasing overhead, only layers that require extra samples are given multisampling in our study.

### v2

@@ Abstract
Crossbar-based in-memory computing (IMC) has emerged as a promising platform for hardware acceleration of deep neural networks (DNNs). However, the energy and latency of IMC systems are dominated by the large overhead of the peripheral analog-to-digital converters (ADCs). To address such ADC bottleneck, here we propose to implement stochastic processing of array-level partial sums (PS) for efficient IMC. Leveraging the probabilistic switching of spin-orbit torque magnetic tunnel junctions, the proposed PS processing eliminates the costly ADC and achieves significant improvement in energy and area efficiency. To mitigate accuracy loss, we develop PS-quantization-aware training that enables backward propagation across stochastic PS. Furthermore, a novel scheme with an inhomogeneous sampling length of the stochastic conversion is proposed. When running ResNet20 on the CIFAR-10 dataset, our architecture-to-algorithm co-design demonstrates up to 16x, 8x, and 10x improvement in energy, latency, and area, respectively, compared to IMC with standard ADC. Our optimized design configuration using inhomogeneous sampling of stochastic PS achieves 130x (24x) improvement in Energy-Delay-Product compared to IMC with full precision ADC (sparse low-bit ADC), while maintaining near-software accuracy at various benchmark classification tasks.
Index Terms:
Deep Neural Networks, Hardware Accelerators, In-Memory Computing, Stochastic Computing1 Introduction
@@
Crossbar-based analog in-memory computing (IMC) has demonstrated great potential for Deep Neural Network (DNN) acceleration by achieving high efficiency and massive parallelism at processing matrix-vector-multiplications (MVMs)[1], which dominates most state-of-the-art DNN workloads[2]. However, the hardware efficiency of IMC sees a severe bottleneck due to the significant overhead of peripheral analog-to-digital converters (ADCs), which are required to ensure robust data communication in a large-scale system (Fig.1). ADC is found to consume over 60-80 of the energy and chip area of IMC hardware [3, 1]. Particularly, the large area overhead of high-precision A-D conversion requires one ADC to be shared by multiple columns in a crossbar. Such design leads to a severe throughput bottleneck since the processing of the array-level outputs across multiple columns has to be done sequentially.
@@
While various recent works have aimed to mitigate the ADC bottleneck [4, 5, 6], minimizing ADC overhead in IMC while still maintaining satisfactory inference accuracy remains a significant challenge. Various factors contribute to the difficulty in addressing the ADC bottleneck. First, since the array-level partial sums (PS) are not represented in DNN algorithms, the standard quantization of activation and weights can not directly address the ADC precision for PS. Second, quantization-aware training with PS quantization desires backward propagation with incorporation of array-level variables, which is challenging as it requires re-designing the computational graph of the DNN model. Third, the required ADC bit precision for MVM may vary significantly depending on both the algorithmic attributes (such as sparsity and DNN layer dimension) and hardware attributes (such as array size and bits per memory cell). Accommodating these varying scenarios requires reconfigurability and flexibility in ADC, leading to increased overhead and design complexity.
@@
In this work, we propose stochastic processing of the array-level partial sum for efficient IMC by leveraging a spin-orbit-torque magnetic tunnel junction (SOT-MTJ) with simple digital circuitry. To address the accuracy drop due to the stochastic PS, a hardware-aware training methodology is developed. Notably, by exploiting the error tolerance of neural networks, our hardware-software co-design achieves the conversion of analog crossbar currents to digital representations with a small number of temporal samplings (1-4 bits for most cases). Such findings address the long-standing issues of large latency incurred from using stochastic bit sequences in computation. Moreover, we conducted extensive system-level hardware simulation to validate that the low-overhead crossbar peripheral based on the proposed spintronic devices/circuits demonstrates significant improvement in energy efficiency and area savings. Our proposed IMC design eliminates the ADC bottleneck and opens up an exciting direction where stochastic computation could play a vital role in designing efficient non-von Neumann DNN accelerators. Our major contributions are the following:
@@
We develop a comprehensive framework of PS quantization-aware training that addresses the accuracy degradation due to the stochastic conversion. In particular, both the device-level stochastic MTJ switching behavior, and key architectural attributes including bit slicing and array size, are incorporated into backward propagation. Our model with 1-bit stochastic PS achieves state-of-the-art accuracy at benchmarking image classification tasks.
@@
We identify that the state-of-the-art quantization of convolutional DNN has severe limitations due to the first convolution layer remaining at high precision. To address this challenge, we develop layer-wise inhomogeneous sampling numbers based on a Monte-Carlo sensitivity analysis to enable aggressive PS quantization on the first convolution layer. As a result, accuracy loss due to quantization is mitigated with minimal additional overhead, demonstrating an advantageous trade-off between efficiency and accuracy.
2 Background
2.1 DNN Acceleration With IMC
@@
Analog IMC with the standard weight-stationary configuration accelerates DNN inference by minimizing the data movement of MVM processing, alleviating the von Neumann memory bottleneck. To process a large-scale DNN workload, large matrices and MVMs with high-precision operands need to be partitioned into multiple crossbar arrays. To ensure robust data communication, an analog crossbar IMC macro is connected with other components through digital-to-analog converters (DAC) at the input and ADC at the crossbar output. The resolution of DAC at each row is typically designed at 1 bit for area saving, while high-precision input activation will be converted to bit streams over multiple time steps (bit streaming). As for high-precision weights, due to the technological limitation in the number of bits per cell, a full-precision weight matrix will be partitioned into several slices of sub-arrays (bit slicing).
@@
The required ADC solution N for array-level PS processing is , where Nrow is the number of activated rows; I is input bits per stream; and W is bits per slice. Since the energy, area, and latency overhead of ADC increase significantly with bit precision [7], ADC becomes the bottleneck of efficiency and performance in IMC design.
2.2 Stochastic Spintronics For MVM in DNN
@@
Several emerging NVM technologies, including resistive memory (ReRAM) and magnetic random access memory (MRAM), have been explored for machine learning acceleration. Recent explorations have demonstrated that spintronic devices exhibit sufficient endurance to realize both synaptic weight storage and neural activation [8]. Particularly, an SOT-MTJ under an excitation current with varying magnitude will have the magnetization switching probability as a sigmoidal function of the current, which provides a direct emulation of a stochastic neuron [9] with significantly higher area/energy efficiency compared to CMOS implementations. Moreover, the separate read/write paths of SOT-MTJs offer immense design flexibility by eliminating the constraint of write current density[10]. All-spin neuro-synaptic processors based on stochastic switching of SOT-MTJ have been proposed [11]. We will explore SOT-MTJs for efficient PS processing of the crossbar output.
2.3 Related Works
@@
Various works have investigated co-designing the IMC hardware architecture and DNN algorithms to reduce the ADC precision. One major thrust focuses on quantization with hardware-aware training. In [12, 13] binary neural networks exhibit improved efficiency for IMC, but the models are limited to 1-bit representation. Recent works [6, 14, 15] implemented bit slicing/streaming to map workloads with multi-bit weights/activations. Aggressively reducing the ADC precision to 1-bit will essentially enable using sense amplifiers. However, 1-bit PS showed sizeable accuracy degradation even after hardware-aware re-training [14, 15]. Moreover, the state-of-the-art quantization-aware training keeps the first convolution layer at full precision [16, 6, 17]. However, the compute-intensive first convolution layer in image processing can dominate the overall computation workloads. Keeping a full-precision first layer severely limits the overall improvement in energy and throughput.
@@
Another major thrust is to exploit sparsity through pruning and re-training to enable lower ADC precision [18, 4]. Higher sparsity reduces the range of possible values of MVM output and thus reduces the ADC resolution requirement. As a result, sparsity-aware IMC with reconfigurable ADC resolution can improve energy and latency. However, it is important to note that the area of the peripheral ADC circuitry remains large in order to handle the highest bit precision of the reconfigurable design. Such large area overhead still hinders the hardware parallelism in IMC architecture.
@@
3 In Memory Computing with Stochastic Partial Sums
The proposed Stox-Net is built on IMC crossbar arrays where current-voltage converters, column-shared MUXs, and peripheral ADCs are replaced by a row of stochastic SOT-MTJs (Fig. 2). We first present the device/circuit design of our stochastic PS processing components and then implement stochasticity-aware training of quantized DNNs with bit slicing considered [17, 16].
3.1 Stochastic PS Processing with MTJ
The stochastic processing based on the current-driven probabilistic switching of SOT-MTJs is shown in Fig. 2. A voltage divider circuit combined with a CMOS inverter functions as an analog-to-stochastic converter using MTJs with tunnel magnetoresistance ratio (TMR) of 350-450% [10]. The switching probability is simulated using a MATLAB-based macro-spin Landau Lifshitz Gilbert (LLG) equation simulator with the SOT current included.
We verify the functionality of the stochastic MTJ-converter by simulating the voltage-divider circuit with MTJs as illustrated in Fig. 2. A modified version of Spintronics Interdisciplinary Center (Spinlib) model [19] is used to model the SOT-MTJ switching. We emulate the crossbar output current using a current source that provides both the positive and negative currents that would accumulate in the crossbar columns (in the range of to ). Note that, in practical operation, the crossbar current is driven by the crossbar input voltages supplied by the digital-to-analog converter (DAC). The terminal T3 of the SOT-MTJ would have a supply voltage that is half of the input supply voltage so that the voltage difference is bi-polar across the heavy metal (HM) layer of the SOT-MTJ, allowing bi-directional current to pass through the HM layer. We demonstrate the operation of the MTJ converter via the waveforms in Fig. 2.
@@ 3.1 Stochastic PS Processing with MTJ
| Supply Voltage | 1 V |
| Ref. MTJ resistance | 140 k |
Our simulations demonstrate a reliable conversion latency of 2ns, indicating rapid response times suitable for high-frequency applications. Energy consumption was measured at for the set operation and at for the reset operation, indicating an average energy consumption of per conversion. The stochastic MTJ converter’s area depicted in Fig. 2 is 0.9108 based on Global Foundries 22FDSOI technology PDK [20]. We scale the area to 28 nm technology node for our study. The dimensions of the reference MTJ are taken as per the GF 22FDX Technology Design Manual, while the dimensions of the SOT MTJ are obtained from Table I. Our proposed SOT-MTJ-based converter is orders of magnitude more energy and area-efficient than ADC, leading to significant improvement of system-level hardware efficiency compared to standard IMC.
3.2 Hardware-aware Training of Quantized DNN
3.2.1 Crossbar Mapping with Stochastic PS
We map large DNN workloads into crossbars with finite array size and limited bits per cell with bit slicing from [1] and input/weight representation from [6]. Following Algorithm 1, mapping a convolution layer of kernel size and input channels using a crossbar array of rows results in PS arrays, where . All input and weight values are then quantized () and split into subarrays (). Weight bits are further split into slices physically mapped to different subarrays, and the input bits are temporally streamed into the crossbar. Each steamed input bit on each sliced subarray will contribute to one PS, leading to partial output that needs to be converted to digital domain and then shift-and-added. SOT-MTJs are connected to process these array-level output currents with the option of multiple samplings. The output of the PS aggregation is then divided by for normalization to [-1, 1] as layer-wise MVM output.
With the binary states of SOT-MTJ encoded as (-1,1), the switching probability versus input current is emulated by the tanh function:
where is the sensitivity parameter. Increasing makes the tanh curve more step-like, approaching a deterministic 1-bit sense amplifier (1b-SA). It has been shown that encoding binary activation to (-1,1) provides better representation capability than encoding to (0,1) [21]. The MTJ’s effective sensitivity in hardware can be altered by tuning the range of crossbar current when mapping MVM operations to hardware.
3.2.2 Backpropagation with Stochastic binarization